If two tangents are drawn from point (p,q)to the parabola y^2=4x such that slope of one tangent is square to other. Then value of p^2+3pq+p

Grade:

1 Answers

Arun
6 years ago
Dear Sanjeev
 
tangent
y = mx + (a/m)
 
Now it passes through p and q
q = mp + (1/m)
 
m^2p - mq + 1 = 0
 
now
m2 = m1^2
 
m1 *m2 = 1/p
 
m1^3 = 1/p
and
m1 + m1^2 = q/p
 
 
Now I think you can solve this.
 
Regards
Arun (askIITians forum expert)